OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation in the interval [0, 2π). Then use a graphing utility to approximate the angle x. (Round your answers to four decimal places.) 4 tan^2 x + 9 tan x − 9 = 0 x = (smallest value) x = x = x = (largest value)

OpenStudy (amistre64):

and what is the quadratic formula?

OpenStudy (anonymous):

ax^2+bx+c=0

OpenStudy (amistre64):

given the format: ax^2 + bx + c = 0 we can derive the quadratic formula, or simply remember it :) \[x = \frac{1}{2a}(-b\pm\sqrt{b^2-4ac})\]

OpenStudy (amistre64):

in this case, we can substitute x, with tan(x) and the rest is the same a = 4, b=9, and c = -9 plug and play

OpenStudy (anonymous):

ok, so i get

OpenStudy (anonymous):

(-9+-√-63))/8

OpenStudy (amistre64):

check your b^2 - 4ac part 81 -4(4*-9) will be positive

OpenStudy (amistre64):

sqrt(225) = 15 (-9+-15)/8 is what i get

OpenStudy (anonymous):

oh yes, you're right

OpenStudy (amistre64):

after that we have to play with trig :)

OpenStudy (anonymous):

ok, this where i get lost a little with what i am trying to answer.

OpenStudy (amistre64):

tan(x) = -24/8 = -3/1 tan(x) = 16/8 = 2 each of those values have 2 angles each that match it

OpenStudy (anonymous):

where are you getting those?

OpenStudy (amistre64):

i figured id use what we found from the quad formula and apply it to solving for tan(x) ... (-9+-15)/8: (-9-15)/8 and (-9+15)/8

OpenStudy (amistre64):

hmmmm, but without the errors :)

OpenStudy (anonymous):

ohhh i see now. ok go on.

OpenStudy (amistre64):

simplify those and the rest is trig-y tan(x) = -3 tan(x) = 3/4 use arctan to find one angle for each one, then add pi to get the other value